OpenStudy (anonymous):

A manufacturing facility pays its employees an average wage of $4.50 an hour with a standard deviation of 50¢. If the wages are normally distributed, what is the percentage of workers getting paid between $3.75 and $5.00 an hour?

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

@rational

OpenStudy (rational):

50 cents is same as 0.5 dollar

OpenStudy (anonymous):

Yes So im looking down the ztable at .5

OpenStudy (rational):

start by finding the zscores for observations 3.75 and 5

OpenStudy (anonymous):

.00009

OpenStudy (rational):

mean = 4.5 standard deviation = 0.5 zscore for observation 3.75 is given by \[\frac{3.75-4.5}{0.5} = -1.5 \]

OpenStudy (rational):

similarly try finding the zscore for obsrevation 5

OpenStudy (anonymous):

.30854

OpenStudy (rational):

how ? show work so that il be able to pinpoint the mistake

OpenStudy (rational):

Nope, don't look at the ztable yet. first we need to find zscores

OpenStudy (rational):

i have worked zscore for the observation 3.75 see if u can work the zscore for the observation 5

OpenStudy (anonymous):

So then it would just be 5-4.5/0.5=x

OpenStudy (rational):

yes what is it

OpenStudy (anonymous):

1?

OpenStudy (rational):

Yes!

OpenStudy (rational):

so we have below so far : zscore for observation 3.75 is ` -1.5` zscore for observation 5 is `1`

OpenStudy (rational):

now open ur ztable, and find the areas corresponding to these zscores

OpenStudy (rational):

subtract both to get the percentage of works getting paid between 3.75 and 5

OpenStudy (rational):

area for the zscore of -1.5 is `0.06681` area for the zscore of 1 is `0.84134` subtract to get the percentage of workers that get paid between 3.75 and 5

OpenStudy (anonymous):

77.4%

OpenStudy (rational):

Yep! good job!

OpenStudy (anonymous):

Awesome thank you!!

OpenStudy (rational):

yw